OpenStudy (18jonea):

Find the radian measures of the angles whose sine is –0.8. Round to the nearest hundredth. –0.93 + 2πn and 4.07 + 2πn –0.93 + 2πn and 5.36 + πn –0.93 + 2πn and 5.36 + 2πn –0.93 + 2πn and 2.21 + 2πn a?

OpenStudy (18jonea):

@robtobey

OpenStudy (18jonea):

@agent0smith

OpenStudy (mysticfury):

whats the question?

OpenStudy (18jonea):

I fixed it

OpenStudy (18jonea):

@agent0smith do you know

OpenStudy (18jonea):

@Mehek14

OpenStudy (18jonea):

@TheSmartOne

OpenStudy (18jonea):

@sweetburger

OpenStudy (18jonea):

@Mehek14

OpenStudy (18jonea):

@jhonyy9

OpenStudy (18jonea):

@Luigi0210

OpenStudy (18jonea):

@paki

OpenStudy (18jonea):

@perl

OpenStudy (18jonea):

@perl

OpenStudy (perl):

You need to solve the equation sin x = -.8

OpenStudy (18jonea):

how would i do that?

OpenStudy (perl):

We can take the inverse sine of both sides. sin^-1 ( sin ( x ) ) = sin^-1( -.8)

OpenStudy (18jonea):

ok so how would i plug thAT into my calc

OpenStudy (perl):

sin^-1 ( sin ( x ) ) = sin^-1( -.8) x = sin^-1( -.8) x = -.927295

OpenStudy (perl):

You can add a multiple of 2pi to this to get coterminal angles.

OpenStudy (18jonea):

ok now what would i do

OpenStudy (perl):

To get the other angle, make use of this identity \( \color{red}{\ sin( \pi - x ) = \sin x }\) Then solve. \( \sin(\pi-x) = -.8 \) \( \pi - x= \sin^{-1}(-.8) \) solve for x \( x = \pi - \sin^{-1}(-.8) \)

OpenStudy (perl):

Do you have a scientific calculator?

OpenStudy (18jonea):

no

OpenStudy (18jonea):

@perl

OpenStudy (18jonea):

@perl

OpenStudy (perl):

I think google has a scientific calculator.

OpenStudy (18jonea):

it wont let me type in "x"

OpenStudy (18jonea):

@perl

OpenStudy (18jonea):

i got 4.07 @perl

OpenStudy (18jonea):

@perl

OpenStudy (perl):

That is correct.